OpenStudy (anonymous):

HELP! Two trains started at the same time from stations which were 360 miles apart and traveled toward each other. The rate of the fast train exceeded the rate of the slow train 10 mph. At the end of two hours, the trains were still 120 miles apart. What was the rate of the train?

OpenStudy (anonymous):

Okay, so I gave you the systems of equations before, but you need to be able to do it yourself. What are good candidates for variables in this problem?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Okay, but what I mean to ask is, what are some unknown values in this word problem. One would be the speed of the slow train... what are some others?

OpenStudy (anonymous):

distance

OpenStudy (anonymous):

Actually, we do know the distance between them.

OpenStudy (anonymous):

oh okay, then the rate

OpenStudy (anonymous):

But we don't know how much distance each one has traveled, so that could be one.

OpenStudy (anonymous):

Sure, so let's have some variables, \(r_1\) for the rate of the fast train and \(r_2\) for the rate of the slow train.

OpenStudy (anonymous):

Now how can we express "The rate of the fast train exceeded the rate of the slow train 10 mph." with the variables \(r_1\) and \(r_2\)?

OpenStudy (anonymous):

\[10 + r2 \]

OpenStudy (anonymous):

We want an equation with both \(r_1\) and \(r_2\)

OpenStudy (anonymous):

But you're very close!

OpenStudy (anonymous):

\[r1 = 10 + r2\]

OpenStudy (anonymous):

Good. Now, let's have \(t\) be the time that has passed. What does \(t\) equal based on: "At the end of two hours"

OpenStudy (anonymous):

\[t= r1+r2\]

OpenStudy (anonymous):

Nope, actually it's just \(t=2\). Alright?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Finally, we have \(d_1\) as the distance the fast train traveled, and \(d_2\) as the distance the slow train traveled. What algebraic equation can we get from: "Two trains started at the same time from stations which were 360 miles " and " At the end of two hours, the trains were still 120 miles apart."

OpenStudy (anonymous):

It won't have \(t\) in it, just \(d_1\) and \(d_2\)

OpenStudy (anonymous):

Hint: remember that \(d_1 + d_2\) is going to be how much closer they got to each other.

OpenStudy (anonymous):

I don't know, I'm a little confused

OpenStudy (anonymous):

is it \[d1 + d2 = 120 ?\]

OpenStudy (anonymous):

Okay, the equation is \(d_1+d_2 = 360 - 120 \)

OpenStudy (anonymous):

Okay, well 360 - 120 is 240

OpenStudy (anonymous):

So the formula we have so far are \[t=2\]\[d_1+d_2 = 240\]\[r_1=r_2+10\]

OpenStudy (anonymous):

5 variables, 3 equations, so we need two more equations... we get them from \(d=rt\). So we have \[d_1=r_1t\]\[d_2=r_2t\]

OpenStudy (anonymous):

Now we have 5 variables, 5 equations. We can solve it!

OpenStudy (anonymous):

Ok, so how do we find d1 and r1?

OpenStudy (anonymous):

okay so we use \(t=2\) to plug \(2\) into every other equation that has \(t\) in it.

OpenStudy (anonymous):

Okay, d1 = r1(2)

OpenStudy (anonymous):

\[d2=r2(2)\]

OpenStudy (anonymous):

so now we can substitute \(2r_2\) into the equation with \(d_2\) in it.

OpenStudy (anonymous):

We have:\[2r_1+2r_2 = 240\]\[r_1=r_2+10\]Certainly you can solve for them now.

OpenStudy (anonymous):

Ok, once again thanks so much, I appreciate it